# Active Learning for Control-Oriented Identification of Nonlinear Systems

### v1

@@ C.2 Proof of correctness for $\texttt{DOED}_{+}$ Algorithm
The proof follows as in the proof of Lemma B.3 of Kakade et al. (2020). Consider a state/input pair $(X_{t},U_{t})$ from rolling out the policy $\pi$ on the system $\phi^{\star}$. Where Kakade et al. (2020) apply the bound on the difference in means between two Gaussian distributions, we use the Gaussian distributions describing the next state update from $X_{t},U_{t}$ under the model described by $\phi$ and by $\phi^{\star}$: $\mathcal{N}(f(X_{t},U_{t};\phi^{\star}),\sigma_{w}^{2})$ and $\mathcal{N}(f(X_{t},U_{t};\phi),\sigma_{w}^{2})$. This results in a bound on the quantity of interest as
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi,\phi}\left[\sum_{t=1}^{T}c(X_{t},U_{t})\right]$
$\leq\sqrt{T\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\left(\sum_{t=1}^{T}c(X_{t},U_{t})\right)^{2}\right]}\sqrt{\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\min\mathopen{}\left\{\frac{1}{\sigma_{w}^{2}}\left\|f(X_{t},U_{t},\phi^{\star})-f(X_{t},U_{t},\phi)\right\|^{2},1\right\}\mathclose{}\right]}.$
Using the fact that the cost is bounded by one, we have
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\left(\sum_{t=1}^{T}c(X_{t},U_{t})\right)^{2}\right]\leq T^{2}$
For the remaining quantity, we may bound
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\min\mathopen{}\left\{\frac{1}{\sigma_{w}^{2}}\left\|f(X_{t},U_{t},\phi^{\star})-f(X_{t},U_{t},\phi\right\|,1\right\}\mathclose{}\right]$ $\leq\frac{1}{\sigma_{w}^{2}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left\|f(X_{t},U_{t},\phi^{\star})-f(X_{t},U_{t},\phi)\right\|^{2}\right]$
$=\frac{1}{\sigma_{w}^{2}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left\|Df(X_{t},U_{t},\tilde{\phi})(\phi^{\star}-\phi)\right\|^{2}\right]$
$\leq\frac{TL_{f}^{2}}{\sigma_{w}^{2}}\left\|\phi^{\star}-\phi\right\|^{2}.$
The equality follows by a first order Taylor expansion of $f$ about $\phi^{\star}$. Here $\tilde{\phi}$ is a convex combination of $\phi$ and $\phi^{\star}$. The final inequality follows from Assumption 2.1. Combining these results concludes the proof. ∎
We now proceed to prove Theorem C.1.
@@
The proof follows by modifying the proof of Theorem 3 in Wagenmaker et al. (2023). In place of the regret minimization algorithm used by Wagenmaker et al. (2023) to estimate the solution to
$\min_{\pi\in\Pi_{\exp}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right],$
we use the policy synthesized directly via optimism over the parameter estimate set $\mathcal{B}$. Let $\phi_{m}$ be the parameter value under which the optimistic policy was synthesized and let $\pi^{m}$ be the optimistic policy. By our assumption that $\phi^{\star}\in\mathcal{B}$, the optimistic policy synthesis provides that
$\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\min_{\pi\in\Pi_{\exp}}\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]\leq\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi^{m},\phi_{m}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right].$
By Lemma C.1, we have that
$\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi^{m},\phi_{m}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]\leq\frac{T^{2}L_{f}}{\sigma_{w}}\left\|\phi^{\star}-\phi^{m}\right\|.$
By the fact that $D(\mathcal{B})$ is the diameter of $\mathcal{B}$, it follows that the above quantity is bounded as
$\operatorname{\mathbb{E}}_{\pi^{m},\phi^{\star}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]-\operatorname{\mathbb{E}}_{\pi^{m},\phi_{m}}\left[\sum_{t=1}^{T}c^{m}(X_{t},U_{t})\right]\leq\frac{T^{2}L_{f}}{\sigma_{w}}D(\mathcal{B}).$
Substituting this bound in place of the bound form the regret minimization step in the proof of Theorem 3 of Wagenmaker et al. (2023), and carrying out the remaining steps from the proof yields that with probability at least $1-\delta$,
$\Phi(\Delta_{\tilde{K}})-\operatorname*{argmin}_{\pi\in\Pi_{\exp}}\Phi(\Sigma^{\pi})\leq\frac{2\beta T^{2}L_{f}^{4}\log(\tilde{K}+1)}{\tilde{K}+1}+B_{\Phi}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log(2\tilde{K}/\delta)}{\tilde{N}}}\right).$
We now consider the correctness of $\texttt{DOED}_{+}$ as it is invoked in 8 by verifying that the conditions to apply the above result are met.
Lemma C.2 (Modified from Corollary 1 of Wagenmaker et al. (2023)).
Let Assumption 2.1 hold. Suppose that $\phi^{\star}\in\mathcal{B}$. Let $D(\mathcal{B})=\sup_{\phi_{1},\phi_{2}\in\mathcal{B}^{-}}\left\|\phi_{1}-\phi_{2}\right\|$. Assume $\mathcal{H}\succ 0$, and define the objective $\Phi(\Lambda)\leftarrow\operatorname{\mathrm{tr}}(\mathcal{H}(\Lambda+\lambda I)^{-1})$ for some $\lambda>0$. Let $B_{\Phi}=TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}/\lambda^{2}$. Then with probability at least $1-\delta$, the data collected at the final round of DOED satisfies
$\Phi\left(\frac{1}{(\tilde{K}+1)\tilde{N}}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
@@
$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(\tilde{N}+1)}{\lambda^{3}(\tilde{N}+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{2\tilde{N}}{\delta}}{\tilde{K}}}\right).$
Proof.
The result follows from Theorem C.1. The smoothness parameters of Assumpton C.1 under the chosen objective are shown in Corollary 1 of Wagenmaker et al. (2023) to be bounded as $B_{\Phi}\leq\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}$ and $\kappa\leq\frac{2T^{2}L_{f}^{4}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{3}}$. ∎
Lemma C.3.
Suppose Assumptions 2.1 and 2.6 hold. Additionally suppose that $\left\|\hat{\phi}-\phi^{\star}\right\|\leq\frac{\lambda_{\min}^{\star}}{4TL_{f}^{2}}$. Then there exists an exploration policy $\pi\in\Pi_{\exp}$ such that
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})\right]\geq\frac{\lambda_{\min}^{\star}}{2}.$
Proof.
The result follows by observing that there exists a policy such that
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right]\geq\lambda_{\min}^{\star}.$
by Assumption 2.6. Then
$\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})\right]-\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})\right]$
$=\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left(Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})-Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right)\right].$
By a first order Taylor expansion of $Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})$ about $\phi^{\star}$, and using the boundedness conditions of Assumption 2.1, we can bound
$\left\|\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}\left(Df(X_{t},U_{t},\hat{\phi})^{\top}Df(X_{t},U_{t},\hat{\phi})-Df(X_{t},U_{t},\phi^{\star})^{\top}Df(X_{t},U_{t},\phi^{\star})\right)\right]\right\|_{\mathsf{op}}\leq 2TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|\leq\frac{\lambda_{\min}^{\star}}{2}.$
Lemma C.4.
Assume that Assumptions 2.1 and 2.6 hold. Additionally, suppose that $\phi^{\star}\in\mathcal{B}$. Let $D(\mathcal{B})=\sup_{\phi_{1},\phi_{2}\in\mathcal{B}}\left\|\phi_{1}-\phi_{2}\right\|$. Assume $\mathcal{H}\succ 0$, and define the objective $\Phi(\Lambda)\leftarrow\operatorname{\mathrm{tr}}(\mathcal{H}(\Lambda+\lambda I)^{-1})$. Let $B_{\Phi}=TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}/\lambda^{2}$. Let $\delta\in(0,\frac{1}{2}]$. Suppose $\lambda\leq\frac{\lambda_{\min}(H)\lambda_{\min}^{\star}}{8\left\|\mathcal{H}\right\|_{\mathsf{op}}d_{\phi}}$, $D(\mathcal{B})\leq\frac{R}{N^{\alpha}}$ and
$N\geq\max\mathopen{}\left\{\left\lceil\frac{4\lambda_{\min}^{\star}T^{4}L_{f}^{8}\log(N+1)}{\lambda^{3}d_{\phi}}\right\rceil^{3},\left\lceil\frac{8\lambda_{\min}^{\star}TL_{f}^{2}\sqrt{2\log(2N/\delta)}}{\lambda^{2}d_{\phi}}\right\rceil^{3},\left(\frac{4T^{3}L_{f}^{3}R\lambda_{\min}^{\star}}{\lambda^{2}\sigma_{w}d_{\phi}}\right)^{1/\alpha}\right\}\mathclose{}$
Then with probability at least $1-\delta$, the data collected at the final round of $\texttt{DOED}_{+}$ satisfy
$\Phi\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,(\tilde{N}\times(\tilde{K}+1))}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right]\right)$ (33)
@@
$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{2\tilde{K}}{\delta}}{\tilde{N}}}\right)$
@@
$\lambda_{\min}\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,((\tilde{K}+1)\times\tilde{N})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)\geq\frac{\lambda_{\min}(\mathcal{H})\lambda_{\min}^{\star}}{4\left\|\mathcal{H}\right\|_{\mathsf{op}}d_{\phi}}.$
Proof.
By Lemma C.2, we have that with probability at leat $1-\delta$,
$\Phi\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,(\tilde{N}\times(\tilde{K}+1))}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)-\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(\tilde{K}+1)}{\lambda^{3}(\tilde{K}+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}D(\mathcal{B})}{\sigma_{w}}+\sqrt{\frac{8\log\frac{2\tilde{K}}{\delta}}{\tilde{N}}}\right)$
$\leq\frac{2T^{4}L_{f}^{8}\left\|\mathcal{H}\right\|_{\mathsf{op}}\log(N+1)}{\lambda^{3}(\lfloor N^{1/3}\rfloor+1)}+\frac{TL_{f}^{2}\left\|\mathcal{H}\right\|_{\mathsf{op}}}{\lambda^{2}}\left(\frac{T^{2}L_{f}R}{N^{\alpha}\sigma_{w}}+\sqrt{\frac{8\log\frac{2N}{\delta}}{\lfloor N^{2/3}\rfloor}}\right)$
$\leq\frac{d_{\phi}}{\lambda_{\min}^{\star}}\left\|\mathcal{H}\right\|_{\mathsf{op}},$
where the final inequality follows from the lower bound on $N$. By Assumption 2.6 paired with Lemma C.3, we have that
$\min_{\pi\in\Pi_{\exp}}\Phi\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]\right)$
$=\operatorname{\mathrm{tr}}\left(\mathcal{H}\left(\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{1},U_{t}^{1};\hat{\phi})\right]+\lambda I\right)^{-1}\right)\leq\frac{d_{\phi}}{\lambda_{\min}^{\star}}\left\|\mathcal{H}\right\|_{\mathsf{op}}.$
We may lower bound
$\Phi\left(\frac{1}{\tilde{N}(\tilde{K}+1)}\sum_{t,n=1}^{T,(\tilde{N}\times(\tilde{K}+1))}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)$
$\geq\frac{\lambda_{\min}\left(\mathcal{H}\right)}{\lambda_{\min}\left(\frac{1}{\tilde{N}\tilde{K}}\sum_{t,n=1}^{T,(\tilde{N}\times\tilde{K})}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right)+\lambda}.$
We conclude by combining these inequalities, and using the assumption that $\lambda\leq\frac{\lambda_{\min}(\mathcal{H})\lambda_{\min}^{\star}}{4\left\|\mathcal{H}\right\|_{\mathsf{op}}d_{\phi}}$. ∎
Lemma C.5.
Consider a policy $\pi\in\Pi_{\exp}$. Suppose $\left\|\hat{\phi}-\phi^{\star}\right\|\leq\min\mathopen{}\left\{\frac{\tilde{\lambda}}{4TL_{f}^{2}},1\right\}\mathclose{}$ for some $\tilde{\lambda}>0$. Define
$\Sigma^{\pi}_{\hat{\phi}}=\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right],$
and suppose that $\Sigma^{\pi}_{\hat{\phi}}\succeq\tilde{\lambda}I$. Then $\Sigma^{\pi}\succeq\frac{\tilde{\lambda}}{2}I$, and
  1. 1.
    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\left(\Sigma^{\pi}\right)^{-1}_{\hat{\phi}}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\tilde{\lambda}}\right)\left\|\hat{\phi}-\phi^{\star}\right\|$
  2. 2.
    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\tilde{\lambda}}\right)\left\|\hat{\phi}-\phi^{\star}\right\|$
Proof.
To begin, we observe that by a first order Taylor expansion of $Df(x,u;\phi^{\star})$ about $\hat{\phi}$, we have that
$\Sigma^{\pi}$ $\succeq\operatorname{\mathbb{E}}_{\pi,\phi^{\star}}\left[\sum_{t=1}^{T}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})^{\top}Df(X_{t}^{n},U_{t}^{n};\hat{\phi})\right]+G,$
where $\left\|G\right\|\leq 2TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|$ by the boundedness assumptions on the derivatives of $f$ from Assumption 2.1. By the assumption that $\left\|\hat{\phi}-\phi^{\star}\right\|\leq\frac{\tilde{\lambda}}{4TL_{f}^{2}}$, it holds that $\left\|G\right\|\leq\frac{\tilde{\lambda}}{2}$. Then it follows that $\lambda_{\min}(\Sigma^{\pi})\geq\lambda_{\min}(\Sigma^{\pi}_{\hat{\phi}})-\left\|G\right\|\geq\frac{\tilde{\lambda}}{2}$.
We proceed to bound $\left\|\left(\Sigma^{\pi}\right)^{-1}-\left(\hat{\Sigma}^{\pi}\right)^{-1}\right\|$. We immediately have that $\left\|\left(\Sigma^{\pi}\right)^{-1}-\left(\hat{\Sigma}^{\pi}\right)^{-1}\right\|\leq\frac{1}{\lambda_{\min}\left(\Sigma^{\pi}\right)\lambda_{\min}\left(\Sigma^{\pi}_{\hat{\phi}}\right)}\left\|\Sigma^{\pi}_{\hat{\phi}}-\Sigma^{\pi}\right\|\leq\frac{2}{\tilde{\lambda}^{2}}\left\|\Sigma^{\pi}_{\hat{\phi}}-\Sigma^{\pi}\right\|$. By Taylor expanding $Df(x,u;\hat{\phi})$ about $\phi^{\star}$, we have from the boundedness assumptions of Assumption 2.1 that $\left\|\Sigma^{\pi}_{\hat{\phi}}-\Sigma^{\pi}\right\|\leq 2TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|+TL_{f}\left\|\hat{\phi}-\phi^{\star}\right\|^{2}\leq 3TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|$. Furthermore, by Lemma B.2, $\left\|\mathcal{H}(\hat{\phi})-\mathcal{H}(\phi^{\star})\right\|\leq\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}\right)\left\|\phi-\phi^{\star}\right\|.$
To conclude, we bound
@@
${\left|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}\right)^{-1}\right)-\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)\right|}$ $\leq\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}\left\|\left(\Sigma^{\pi}\right)^{-1}-\left(\hat{\Sigma}^{\pi}\right)^{-1}\right\|\leq\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}\frac{6TL_{f}^{2}\left\|\hat{\phi}-\phi^{\star}\right\|}{\tilde{\lambda}^{2}},$
and
${\left|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)-\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\left(\Sigma^{\pi}_{\hat{\phi}}\right)^{-1}\right)\right|}$ $\leq\frac{d_{\phi}}{\tilde{\lambda}}\left\|\mathcal{H}(\hat{\phi})-\mathcal{H}(\phi^{\star})\right\|$

### v2

@@ Appendix B Results for smooth nonlinear systems
Express
$\nabla_{\phi}v^{\top}\mathbf{FI}^{\pi}(\phi)v=\nabla_{\phi}\int\sum_{t=1}^{T}v^{\top}Df(X_{t},\pi(X_{t});\phi)Df(X_{t},\pi(X_{t});\phi)^{\top}vp_{\phi}(X_{1:T+1})d_{X_{2:T+1}},$
where $X_{1}=0$ and $p_{\phi}(X_{1:T+1})=\Pi_{t=1}^{T}p_{w}(X_{t+1}-f(X_{t},\pi(X_{1:t});\phi)).$
The gradient is then given by
$\nabla_{\phi}v^{\top}\mathbf{FI}^{\pi}(\phi)v$ $=\int\left(\nabla_{\phi}\sum_{t=1}^{T}v^{\top}Df(X_{t},\pi(X_{1:t});\phi)Df(X_{t},\pi(X_{1:t});\phi)^{\top}v\right)p_{\phi}(X_{1:T+1})d_{X_{2:T+1}}$
$+\int\sum_{t=1}^{T}v^{\top}Df(X_{t},\pi(X_{1:t});\phi)Df(X_{t},\pi(X_{1:t});\phi)^{\top}v\left(\nabla_{\phi}\log\left(p_{\phi}(X_{1:T+1})\right)\right)p_{\phi}(X_{1:T+1})d_{X_{2:T+1}}.$
By Assumption 2.1, the first term is differentiable, and the norm is bounded by $2TL_{f}^{2}$. For the second term, observe that
$\nabla_{\phi}\log\left(p_{\phi}(X_{1:T+1})\right)=\nabla_{\phi}\left(-\frac{1}{2\sigma_{w}^{2}}\sum_{t=1}^{T}\left\|X_{t+1}-f(X_{t},\pi(X_{1:T});\phi)\right\|^{2}\right)=\frac{1}{\sigma_{w}^{2}}\sum_{t=1}^{T}D_{\phi}f(X_{t},\pi(X_{1:t};\phi)W_{t}.$
Therefore, the norm of the second term may be bounded by $\frac{T^{2}L_{f}^{3}\sqrt{d_{\mathsf{X}}}}{\sigma_{w}}$ by appealing to the triangle inequality, submultiplicativity, the bound $\mathbf{E}[\left\|g_{t}\right\|]\leq\sqrt{n}$ if $g_{t}$ is a standard multi-variate random vector of dimension $n$.
Appendix C Proof of Main Excess Cost Bound, Theorem 3.2
First we present a result stating that using the estimated parameter in place of the true parameter leads to an adequate approximation for the ideal exploration objective.
Lemma C.1.
Consider a policy $\pi\in\Pi_{\exp}$. Suppose $\left\|\hat{\phi}-\phi^{\star}\right\|\leq\frac{\tilde{\lambda}\min\mathopen{}\left\{\sigma_{w},1\right\}\mathclose{}}{6T^{2}\max\mathopen{}\left\{1,L_{f}\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}}$ for some $\tilde{\lambda}>0$. Assume that $\mathbf{FI}^{\pi}(\hat{\phi})\succeq\tilde{\lambda}I$. Then $\mathbf{FI}^{\pi}(\phi^{\star})\succeq\frac{\tilde{\lambda}}{2}I$, and
  1. 1.
    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right)\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\tilde{\lambda}}\right)\left\|\hat{\phi}-\phi^{\star}\right\|$
  2. 2.
    $\operatorname{\mathrm{tr}}\left(\mathcal{H}(\hat{\phi})\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right)\leq\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right)+\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}},d_{\phi},\left\|\mathcal{H}(\phi^{\star})\right\|,\frac{1}{\tilde{\lambda}}\right)\left\|\hat{\phi}-\phi^{\star}\right\|$
@@
Begin by bounding $\left\|\mathbf{FI}^{\pi}(\hat{\phi})-\mathbf{FI}^{\pi}(\phi^{\star})\right\|=\sup_{v\in\mathbb{R}^{d_{\theta}},\left\|v\right\|=1}v^{\top}\mathbf{FI}^{\pi}(\hat{\phi})v-v^{\top}\mathbf{FI}^{\pi}(\phi^{\star})v$. By a first order Taylor expansion, it holds that for any unit vector $v\in\mathbb{R}^{d_{\theta}}$,
$v^{\top}\mathbf{FI}^{\pi}(\hat{\phi})v-v^{\top}\mathbf{FI}^{\pi}(\phi^{\star})v\leq\left\|\nabla_{\phi}v^{\top}\mathbf{FI}^{\pi}(\tilde{\phi})v\right\|\left\|\hat{\phi}-\phi^{\star}\right\|.$
for some $\tilde{\phi}=\alpha\hat{\phi}+(1-\alpha)\phi^{\star}$, $\alpha\in[0,1]$. By Lemma B.4, it holds that
@@
$\left\|\nabla_{\phi}v^{\top}\mathbf{FI}^{\pi}(\tilde{\phi})v\right\|\leq 3\frac{T^{2}\max\mathopen{}\left\{L_{f},1\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}}{\min\mathopen{}\left\{1,\sigma_{w}\right\}\mathclose{}}.$
Then by the assumed bound on $\left\|\hat{\phi}-\phi^{\star}\right\|$, we find that $\left\|\mathbf{FI}^{\pi}(\hat{\phi})-\mathbf{FI}^{\pi}(\phi^{\star})\right\|\leq\frac{\tilde{\lambda}}{2}$. Therefore, $\mathbf{FI}^{\pi}(\phi_{\star})\succeq\frac{\tilde{\lambda}}{2}$.
We also have that
@@
$\left\|\mathbf{FI}^{\pi}(\hat{\phi})^{-1}-\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right\|\leq\frac{1}{\lambda_{\min}\left(\mathbf{FI}^{\pi}(\hat{\phi})\right)\lambda_{\min}\left(\mathbf{FI}^{\pi}(\phi^{\star})\right)}\left\|\mathbf{FI}^{\pi}(\hat{\phi})-\mathbf{FI}^{\pi}(\phi^{\star})\right\|\leq\frac{6T^{2}\max\mathopen{}\left\{L_{f},1\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}}{\tilde{\lambda}^{2}\min\mathopen{}\left\{1,\sigma_{w}\right\}\mathclose{}}\left\|\hat{\phi}-\phi^{\star}\right\|$
Furthermore, by Lemma B.2, $\left\|\mathcal{H}(\hat{\phi})-\mathcal{H}(\phi^{\star})\right\|\leq\mathsf{poly}\left(L_{\pi^{\star}},L_{f},L_{\theta},L_{\mathsf{cost}},\sigma_{w}^{-1},T,d_{\mathsf{X}}\right)\left\|\phi-\phi^{\star}\right\|.$
To conclude, we bound
${\left|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\phi^{\star})^{-1}\right)-\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right)\right|}$ $\leq\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}\left\|\mathbf{FI}^{\pi}(\phi^{\star})^{-1}-\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right\|$
$\leq\left\|\mathcal{H}(\phi^{\star})\right\|d_{\phi}\frac{6T^{2}\max\mathopen{}\left\{L_{f},1\right\}\mathclose{}^{3}\sqrt{d_{\mathsf{X}}}}{\tilde{\lambda}^{2}\min\mathopen{}\left\{1,\sigma_{w}\right\}\mathclose{}}\left\|\hat{\phi}-\phi^{\star}\right\|$
@@
${\left|\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right)-\operatorname{\mathrm{tr}}\left(\mathcal{H}(\phi^{\star})\mathbf{FI}^{\pi}(\hat{\phi})^{-1}\right)\right|}$ $\leq\frac{d_{\phi}}{\tilde{\lambda}}\left\|\mathcal{H}(\hat{\phi})-\mathcal{H}(\phi^{\star})\right\|$